# Anatomy of an ionized bubble: NIRCam grism spectroscopy of the double-peaked Lyman- emitter COLA1 and its environment

### v2

@@ 4.5 Escape fraction of Ly and LyC photons
Empirically, the LyC escape fraction is correlated with the peak separation of the Ly line (e.g., Izotov et al., 2018). The peak separation of km s-1 observed in the VLT/X-shooter spectrum suggests (see Matthee et al., 2018). Furthermore, Naidu et al. (2020) used a bayesian inference model to fit a power-law relation between the LyC escape fraction and the star formation rate surface density, , from observational constrains of the reionization timeline and the galaxy size evolution. Following this relation, we obtain (LyC) . Lastly, Chisholm et al. (2022) found a relation between the UV slope and the LyC escape fraction, for a sample of SFGs at with up to . We extrapolate their best fit to get (LyC) , hence a lower limit of (LyC) .

### v3

@@ 4.5 Escape fraction of Ly and LyC photons
Empirically, the LyC escape fraction is correlated with the peak separation of the Ly line (e.g., Izotov et al., 2018). The peak separation of km s-1 observed in the VLT/X-shooter spectrum suggests (see Matthee et al., 2018). Furthermore, Naidu et al. (2020) used a bayesian inference model to fit a power-law relation between the LyC escape fraction and the star formation rate surface density, , from observational constrains of the reionization timeline and the galaxy size evolution. Following this relation, we obtain a lower limit of (LyC) . Lastly, Chisholm et al. (2022) found a relation between the UV slope and the LyC escape fraction, for a sample of SFGs at with up to . We extrapolate their best fit to get (LyC) , hence a lower limit of (LyC) .